OpenStudy (anonymous):

A recipe that makes 3 servings calls for 3/5 cup of apple juice and 3/4 cup of cranberry juice. How many cups of juice are there per serving?

OpenStudy (anonymous):

A. 4 1/20 B. 1 7/20 C. 2/3 D. 9/20

OpenStudy (johnweldon1993):

Alright so together we have \[\large \frac{3}{5} + \frac{3}{4} = \frac{27}{20}\] per 3 servings...so we can just divide that by 3 to get the per 1 serving... \[\large \frac{27}{20} \div 3\] we flip the second number and change this to multiplication \[\large \frac{27}{20} \times \frac{1}{3} = \frac{27}{60} = \frac{9}{20}\]

OpenStudy (anonymous):

how did u get 27/20

OpenStudy (johnweldon1993):

Well we have 2 types of juices going into this recipe...so since they want to know how much juice (that would be a mixture of the 2) so we need to add our 2 fractions...

OpenStudy (anonymous):

yeah but the adding is wrong

OpenStudy (johnweldon1993):

So \[\large \frac{3}{5} + \frac{3}{4}\] we need first find a common denominator..a number both 5 and 4 go into evenly..20 seems about right so we need to multiply the whole first fraction by 4 to make the denominator 20...and the whole second fraction by 5 to make that denominator 20 so \[\large \frac{4}{4} \times \frac{3}{5} + \frac{5}{5} \times \frac{3}{4}\] this will come out to \[\large \frac{12}{20} + \frac{15}{20}\] which will equal \[\large \frac{27}{20}\]

OpenStudy (johnweldon1993):

Dont worry about it, but from there does everything else make sense?

OpenStudy (anonymous):

sorry sorry i forgot the denominator and yes it dose make sense